My wallet contains $3409.00 in ones, twenties, and fifties. Find the number of each kind of bill if there are 6 fewer ones than twenties and 24 more twenties than fifties.
Answer by Paul(988)

ones: x+24-6
tewnties: x+24
Fifties: x
50(x)+(x+24-6)+20(x+24)=3409
50x+x+24-6+20x+480=3409
71x+498=3407
71x=2911
x=41
So therefore there where 41 fifties.